### v1

# Customizing Motion in Text-to-Video Diffusion Models

Abstract
11footnotetext: *Work done in part as an intern with Adobe Research.11footnotetext: Czech Institute of Informatics, Robotics and Cybernetics at the Czech Technical University in Prague.We introduce an approach for augmenting text-to-video generation models with customized motions, extending their capabilities beyond the motions depicted in the original training data. By leveraging a few video samples demonstrating specific movements as input, our method learns and generalizes the input motion patterns for diverse, text-specified scenarios. Our contributions are threefold. First, to achieve our results, we finetune an existing text-to-video model to learn a novel mapping between the depicted motion in the input examples to a new unique token. To avoid overfitting to the new custom motion, we introduce an approach for regularization over videos. Second, by leveraging the motion priors in a pretrained model, our method can produce novel videos featuring multiple people doing the custom motion, and can invoke the motion in combination with other motions. Furthermore, our approach extends to the multimodal customization of motion and appearance of individualized subjects, enabling the generation of videos featuring unique characters and distinct motions. Third, to validate our method, we introduce an approach for quantitatively evaluating the learned custom motion and perform a systematic ablation study. We show that our method significantly outperforms prior appearance-based customization approaches when extended to the motion customization task.
1 Introduction
Recent advancements in text-to-video synthesis have yielded remarkable outcomes [2, 43, 40, 18, 42, 50, 28], marking progress in a traditionally challenging area of video generation. However, the capabilities of these models remain confined to their training data. Conveying specific motions not covered in the training set is challenging. For example, consider the examples of the depicted ‘Carlton’ dance shown in Figure 1 (left). To describe the ‘Carlton’ dance in a text prompt, it is difficult to describe with natural language the nuances of the swinging arms and the timings of the poses in the dance. The ability to generate a new motion unknown to the model for one or more desired characters with the flexibility of using text to control the subjects’ appearance, as shown in Figure 1 (right), can be particularly useful for animation or content creation applications. This ability leads us to a question: How can we harness an existing text-to-video model’s knowledge about motion and appearance while augmenting them with new, custom motions?
While appearance-based customization techniques in text-to-image synthesis have enabled tailored image generation [9, 22, 32],their scope is confined to still images. This setting introduces a limitation where only object customization is feasible. Recent works have extended a pretrained text-to-image synthesis model with trainable temporal layers and applied one-shot tuning over a single video [7, 45]. While these approaches enable text-based stylization of the single video, the generated videos have motions that adhere closely to the timing and spatial layout of the input video and cannot generalize to any other motions. As a result, for an input video’s depicted motion, these methods cannot combine it with other motions, apply it to multiple people in a scene, or vary its viewpoint/timing.
In this work, we present a new approach for customizing text-to-video diffusion models to a novel motion regardless of the appearance of the motion’s subject. Given a few videos depicting a new motion, e.g., dance or gesture, our goal is to map the novel motion to a unique text token (“V*”). Given the unique text token, we can generate videos with varied subjects undertaking the novel motion (cf., Figure 1 right).
As our first contribution, our approach goes beyond prior work by leveraging the rich motion and appearance prior information contained in a pretrained text-to-video generation model. We augment this model with new custom motions by fine tuning the temporal layers and spatial cross-attention keys/values, which we show are crucial for modeling the novel motion. To maintain the model’s motion prior and avoid overfitting to the new motion, we introduce a new approach for regularization tailored to videos.
As our second contribution, we show that our approach learns to generalize and hence we can invoke the custom motion in novel applications. For example, we first show how to apply a custom motion to multiple people in a scene. Second, we show how to invoke the newly learned custom motion in combination with other motions. Finally, we show how to perform multimodal customization over motion and appearance.
As our third contribution, to validate our method, we introduce a new approach for quantitatively evaluating the fidelity of the generated custom motions along with the model’s ability to respect the subject’s desired appearance as described in a text prompt. We use this evaluation protocol to perform a systematic ablation of our method and find that we significantly outperform baseline appearance-based customization approaches when extended to the motion customization task.
2 Related Work
Text-to-image customization. Previous approaches to customization in text-to-image generation models are focused on introducing a novel subject or object [32, 22, 9, 6, 10, 44, 16, 37, 34, 47, 24, 46, 33, 5, 26, 51]. One strategy fine-tunes a subset of the parameters of the generation network given a few images of a subject and its descriptions [32, 22]. Through this optimization, a new text-image mapping is created between the subject depicted in the images and the textual description. At test time, the subject is invoked via the description but can be now used in a free-form context. Similarly, Gal et al. [9] showed that this mapping can be created purely by optimizing a text embedding of the textual description without modifying the generation network, at a price of image quality. More recent approaches investigate efficient formulations for modifying an approximation of the parameters responsible for the text-image mapping in the generation network [16, 37]. Other methods, alternatively to adjusting the model’s parameters, introduce an image encoder that can be directly plugged into the pre-trained network [10, 44, 51]. Alternatively to customizing individual concepts, Huang et al. [20] propose to learn relations between two subjects from a set of images via a similar technique to Gal et al. [9]. The text-to-image customization is limited to static objects due to the static nature of still images. In contrast, we focus on customization with respect to a given motion depicted in a small set of videos.
Controllable video generation. The study of controllable video generation is widely researched; we summarize briefly here works that are closest to ours. Han et al. [15] use multi-modal input like sketches or semantic layout to control the synthesized content. Previous works learn GAN representations that disentangle motion and structure [39] or decompose foreground and background to manipulate the forground only [19]. Chan et al. [4] use a single source video of a dancing person to generate a novel video depicting the motion transferred to a new subject. Geyer et al. [13] propose a method for semantic video editing of real-world videos, using a text-to-image model. Our approach differs from these prior works in that we learn a generalized motion from a few video examples. Then, the motion can be used in new contexts with text conditioning, and the motion can be used in an array of new applications where text controls for the number of subjects performing the motion, their appearance, and the background. While previous works focus primarily on transferring motion between humans, our approach generalizes the motion pattern such that it can be applied also to non-human subjects.
Video synthesis with image generation models. Prior work has used text-to-image generation models for video generation. Guo et al. [14] learn a motion prior network with text-video data and leverage the prior in personalized text-image models. Similarly, prior work has shown video generation with text-to-image models [35, 48, 21]. Ge et al. [12] demonstrates an approach for zero-shot text-to-video generation by designing a video noise prior. Our work differs in that we use a video generation model directly to learn new motions. Wu et al. [45] inverts an existing video frame-by-frame using a pre-trained text-to-image model then, using the inverted noise maps and text guidance, synthesizes a novel video. This method edits an existing video, while our focus is on generating completely new videos from scratch.
Text-to-video models. Recently, several methods have demonstrated text-to-video generation [2, 43, 40, 18, 42, 50, 28]. An important line of research involves studying architectures and optimization for training video generation models. A high-quality model is a prerequisite for our method. In this work, we use the model from Wang et al. [42]. Molad et al. [28] show qualitative results of adding a custom subject to the base video model. They, however, do not explore customization beyond static objects [28]. The idea of multi-modal customization, however, facilitates user-controllability and creative applications, since no matter how powerful the model, there are always novel subjects and motions that cannot be all contained in the training set.
3 Approach
We aim to customize text-to-video models with respect to a novel motion. That is, we wish to introduce novel dynamic concepts in the model’s knowledge base and associate them with unique text tokens. In this section, we first give a preliminary description of text-to-video diffusion models, followed by our approach for motion customization.
3.1 Text-to-Video Diffusion Model Preliminaries
Diffusion models are probabilistic models that can approximate distributions through a gradual denoising process [17, 36]. Given a Gaussian noise sample, the model learns to iteratively remove a portion of noise until a sample is within the approximated distribution. The objective function of this process minimizes the distance of the predicted noise and the sampled noise. Latent Diffusion Models (LDMs) [31] are diffusion models operating in the latent space instead of pixel space. Instead of learning the denoising process in a complex high-dimensional space, the videos are first encoded into a lower-dimensional vector through a learned encoder-decoder model.
Specifically, a video is represented with a latent vector, and a text condition is encoded through a text-embedding from a pre-trained text encoder. The initial noise map is sampled from a Gaussian distribution . For the diffusion timestep sampled from a probability density , the noisy latent sample can be computed from the noise prediction by , where is the original video and controls the amount of noise added at each diffusion timestep according to a noise scheduler [36, 17]. The model with parameters is trained with the following weighted denoising loss,
where is a user-defined variable that controls the sample quality. At inference time, a new video can be generated by sampling a Gaussian noise sample and a text prompt, denoising them with the learned model.
3.2 Approach for Motion Customization
We illustrate our overall approach for motion customization in Figure 2. Let the motion be represented through a small exemplar set of videos and corresponding text prompts . The motion can be performed by different subjects across different backgrounds, and the commonality within the videos is purely the dynamic movement. We choose a generic textual description across the videos, such as “a person doing the V*”. In practice, we select unique tokens like “pll” as “V*”. To customize a text-to-video model’s parameters , we fine-tune the model’s parameters by minimizing the diffusion objective summed over the exemplar set ,
At test time we can generate novel videos of the target with variations controlled by text. Next, we describe the important components of our motion customization approach.
Choice of customization parameters. The quality and generalization ability of the novel motion depends on the choice of the model parameters updated during customization. A text-to-video diffusion model has parameters that can be categorized to those operating on the temporal dimensions and those operating on the spatial dimensions (per-frame) . Let be the keys and values parameters of the spatial cross-attention modules.
The temporal layers are transformers and temporal convolutions, and are responsible for modelling temporal patterns across frames. In Section 4, we empirically show that the temporal layers alone do not effectively model a new motion pattern due to time-varying appearance changes. For example, consider an object rotating in 3D, which requires the model to generate the appearance of disoccluded surfaces.
To faithfully learn a motion pattern, we also modify a subset of parameters in the spatial layers of the model. As illustrated in Figure 2, our approach fine tunes the spatial keys/values and temporal parameters. Note that in image customization [22, 38] and model editing [11], it has been shown that the spatial keys and values of the cross-attention modules in text-to-image models are sufficiently powerful to represent the appearance of new concepts. Our experiments also show the effectiveness of adjusting those parameters for the motion customization task.
Video regularization. Prior work has shown that directly optimizing Equation (2) leads to forgetting related concepts or the concept category [32, 25, 23]. For example, if the concept is a specific person, all people start resembling that person. To mitigate this issue, prior work has utilized a regularization set that ensures that this knowledge is preserved. For example, Ruiz et al. [32] proposed collecting a regularization set via generation from the original model. Kumari et al. [22] proposed using pairs of real images and text.
In contrast to prior work, we seek to mitigate forgetting of related motions that the model has already learned. To address this goal, we consider a video-based regularization. Let be a regularization set of paired videos with text descriptions that have similar but not identical motion to the target custom motion videos . For example, when learning the ‘Carlton’ dance, we chose a regularization set of real videos containing people dancing. For customization, we optimize the diffusion objective over the target custom videos and regularization set :
Empirically, we show in Section 4 that using real videos for the regularization set is superior to using generated videos from the model. We find that using generated videos degrades the quality of the customized model substantially.
Learning the motion patterns. To facilitate learning the common motion in , during the training process, we aim to put emphasis on the motion pattern of the videos, rather than the appearance or subject presenting the motion. For example, in the case of the ‘Carlton’ dance in Figure 1 we wish to capture the motion pattern of the dance rather than the appearance of the background or the individual performers. The denoising process in diffusion models samples a Gaussian noise and then gradually removes the noise. The initial noise as well as early denoising steps have large influence on the output overall dynamic structure of the video, whereas the later stages correspond to finer details [41]. To focus on the dynamic structure of the videos and de-emphasize the appearance of the subjects performing the motion, we define a timestep sampling strategy. In particular, we build on [20] who develop a non-uniform sampling strategy for generating still images. Here we adopt it for generating dynamic video content. In detail, instead of uniformly sampling the denoising steps in Equation (1), we define a probability distribution over the timesteps that focuses more on the earlier denoising steps and hence emphasizes learning the overall dynamic motion pattern rather than the fine appearance details of the background or the individual subjects. The parameter increases the skewness of the function. We use for all of our experiments.
Customizing appearance and motion. Our method can be easily extended to customizing both appearance and motion. The appearance of a new subject is represented through an exemplar dataset of pairs of still images and captions describing the images. To learn the appearance of a person we follow [22] and update the spatial layers of the model, and optimize the text embedding associated with our subject. We customize the model first with the motion, then train both the appearance and the subject jointly, choosing an example of either dataset in the mini-batch. When learning the subject we treat the images as a one-frame video and only update the selected spatial layers.
| Motion accuracy | Text Alignment | |
| Textual Inversion [9] | 0.3 | 0.2733 |
| Custom Diffusion [22] | 10.5 | 0.2788 |
| Dreambooth [32] | 28.4 | 0.2796 |
| Ours | 70.6 | 0.2818 |
| Spatial layers | Accuracy | Copy |
| None | 62.3 | 5.9 |
| All | 68.7 | 17.0 |
| K,V (Ours) | 70.6 | 8.7 |
| Trans. layers | Accuracy | Copy |
| None | 22.7 | 1.4 |
| All (Ours) | 70.6 | 8.7 |
| Text Token | Accuracy | Copy |
| ✓ | 75.5 | 22.7 |
| ✗ (Ours) | 70.6 | 8.7 |
| Sampling | Accuracy | Copy |
| Uniform | 66.9 | 15.4 |
| Coarse-noise (Ours) | 70.6 | 8.7 |
| Prior | Accuracy | Copy |
| None | 43.9 | 1.2 |
| Synthetic | 48.3 | 3.7 |
| Real (Ours) | 70.6 | 8.7 |
4 Experiments
In this section, we quantitatively evaluate our approach (Section 4.1), perform an ablation study (Section 4.2), and show qualitative results (Section 4.3).
Implementation details. In all our experiments we use the ZeroScope text-to-video diffusion model [3] as the pre-trained network. Our implementation is building on the public repository [8]. In our concept exemplar dataset , we apply spatial and color augmentation on the video examples. We train the model with a mini-batch size 1 on a single A100 Nvidia GPU. We train the models up to 3000 training steps with learning rate .
4.1 Quantitative Evaluation
We quantitatively evaluate our approach by computing metrics corresponding to the quality of the generated motion and the overall fidelity of the generated video with respect to the input text prompt.
Dataset. We quantitatively evaluate our approach on the Jester Dataset [27]. The dataset contains 148,092 crowd-sourced videos of 27 human gestures by diverse actors performing a gesture in front of a static camera. We select 5 gesture classes as our target motions – Swiping Down, Drumming Finders, Rolling Hand Forward, Shaking Hand, Sliding Two Fingers Up. From each class, we choose at random 10 videos from the validation set as our training set. To learn the motion, we caption the videos with a generic text description “a person doing a V* sign”. As a regularization set for the motion learning, we also choose 50 videos from each of the dataset classes Doing Other Things, No gesture and assign them text captions “a person doing other things” and “a person in front of a computer”. We design a test set containing 100 text prompts that detail a person’s appearance, for example “A female firefighter doing the V* gesture” or “A Japanese toddler doing the V gesture”. We fix 3 random seeds per text prompt to evaluate our experiments.
Motion accuracy metric. The Jester dataset has been widely adopted for the task of action recognition. We leverage this for our motion accuracy metric, measuring gesture recognition accuracy with a video classifier, pre-trained on the Jester training set [49]. The classifier achieves 94.78% accuracy across all gesture classes, making it suitable for our evaluation. Note that the simple test set prompts typically result in generated videos that adhere to the spatial layout of the gesture recognition classifier’s training videos with a single person, making them in-distribution.
Text alignment metrics. To regulate overfitting, we measure how faithfully the model adheres to the appearance described in the text prompt with an appearance score defined as the mean CLIP score between the generated frames and the part of the text prompt describing the person’s appearance (e.g., “A female firefighter”). We empirically find that the appearance score is a good proxy to determine when the model starts to overfit. For a fair comparison across all models, we fine-tune each model until this appearance score reaches a chosen value (here 0.265), determined empirically based on the visual quality of the results. Please refer to the Supplemental Material for an additional discussion on the tradeoff between motion accuracy and appearance alignment.
To assess how well the generated video corresponds to the overall text prompt (both motion and appearance), we measure the text alignment score as a mean CLIP score between the video frames and a full caption written in natural language. For example, for the class Swiping Down and a text prompt “A female firefighter doing the V* gesture”, a full caption would be “A female firefighter swiping hand down.” We use this score as another criterion to compare different methods.
Comparison with image customization methods. We select three well-established image customization approaches: Textual Inversion [9], Dreambooth [32] and Custom Diffusion [22], and adapt them to the motion customization task. Textual Inversion [9] optimizes a text token embedding to learn a novel subject. For Dreambooth [32], we optimize all spatial layers of the model and use a synthetic video regularization set following their prior preservation loss. Custom Diffusion trains the key and value projection matrices in the cross-attention layers of text-to-image diffusion models, optimizes a text token embedding, and uses a regularization set containing real videos. We report the motion accuracy and text alignment metrics in Table 1 and show qualitative comparison in Figure 3. We observe that Textual Inversion completely fails to learn the motion. We hypothesize that this failure is because the text encoder has been trained only on image-text pairs, and the embedding fails to generalize to the video domain. Additionally, the text embedding is injected only into spatial layers in the cross-attention layers, and not the temporal ones. Alternatively, because we are learning an unseen motion pattern, which is more complex than a static object, the embedding does not have enough parameters to learn it. On the other hand, Dreambooth and Custom Diffusion learn to adhere to the spatial structure of the training videos and some hand manipulation. However they fail to accurately reproduce the motion and produce a temporally smooth video. In contrast, our approach yields over a 2 improvement in motion accuracy and more faithfully adheres to the text prompt.
4.2 Ablation study
We perform an ablation study of our method’s components and report the motion accuracy metric in Table 2. In addition, we report a copying score (“Copy”) that measures how much of the training data appearance leaks into the generated data. We use the SSCD description score [29] to detect image copying between the training videos and the generated videos. We compute our copying score as the maximum SSCD description score between the generated frames of the video and first frames of the training examples, and report the percentage of samples that are above a specified threshold. For our experiments, we manually determine a threshold of 0.25, which corresponds to significant copying of the appearance of the person or the background (please see examples in the Supp. Mat.).
First, we study the choice of the optimized parameters. As in the case of the image customization methods, it is clear that optimizing the spatial layers of the model is not enough to learn a novel motion pattern, as shown in Table 1(b). We can further observe that training the key and value projections in the cross-attention layers of the spatial layers achieves a good balance in terms of accuracy and copying score, as shown Table 1(a). We observe a two-fold reduction in the copying score when training only those parameters and an 8% improvement in motion accuracy compared to not training any spatial parameters.
Next, in Table 1(c), we consider the role of optimizing the text token embedding. We notice that when jointly training the text token embedding with the model’s parameters, the motion accuracy is higher than when training only the model parameters. However, it also leads to nearly three times as much memorization of the training video appearance, as evidenced by the significantly increased Copy score. We hypothesize that this indicates that the text token embedding is learning something related to the appearance of the training videos.
Consequently, we test our sampling strategy. We compare a model trained with our motion pattern sampling strategy (“Coarse-noise”) to a model that simply samples diffusion time steps uniformly (“Uniform”). As shown in Table 1(d), our sampling strategy improves the motion accuracy and reduces the appearance copying. Following our intuition regarding learning motion patterns rather than fine-grained details, our model is less prone to replicating elements from the training examples.
Lastly, we study the effect of the regularization prior in Table 1(e). We train three models with the same design choice as our method, yet we choose the regularization set to be either: (i) empty, (ii) a synthetic set of videos generated from the original text-to-video model, or (iii) containing real videos. We observe a significant drop in performance both without using a prior as well as using a synthetic prior. Interestingly, despite all settings using the same fine-tuning set, a high-quality regularization set containing real videos improves the overall training. Overall, our method achieves a good balance between high motion accuracy and low copying score.
4.3 Qualitative experiments
To further demonstrate the effectiveness of our approach, we choose the motions: Dab, Airquotes, and 3D rotation, which the original model does not know or cannot reliably perform. For the first two motions, we curate videos from the internet searching for the description of the motion. We use between 5 and 6 videos for each motion and caption the videos with a generic text description such as “A person doing the V* sign” or “A person V*”. To learn 3D rotation, we select 10 sequences of distinct objects from the CO3D Dataset [30] and caption the videos with the text prompt “A camera V* around an [object]”, where [object] is an object class. The results are shown in Figure 4. For visualization, we query for objects not present in the training videos. For the regularization set, we select 100 video and text pairs from the WebVid-10M Dataset [1]. We fine-tune the models for 2000 steps. Please see our companion webpage for additional results and comparison with the pre-trained model.
Motion and appearance customization. We combine the appearance of a specific person with a gesture motion Drumming Fingers from the Jester Dataset. We use the publicly available images of a particular person from the CustomConcept101 Dataset [22], and annotate them with generic text captions {“A X* man”, “A close-up of a X* man”, “A X* man smiling.”}. For the motion dataset, we choose the same 10 videos as in Section 4.1 and a generic text description “a person doing a V* sign”. Our qualitative results are shown in Fig. 6.
Comparison with a reference-based method. We compare our method with Tune-A-Video [45], a video editing method that, given a source video, inflates a text-to-image model to learn a motion pattern from a single video. This method allows generation of novel videos with the same pattern but varying appearance. We adapt Tune-A-Video to fine-tune from multiple videos using the same training set as our setup. In Figure 5 we compare the two methods. We test the generalization abilities of the two methods, firstly combining the novel motion with a different motion: “doing the V* gesture while eating a burger with other hand”. We can see that our method can generate the two motions. Next, we test the motion variation. We show that we can manipulate the execution of the motion with a natural language description; here we specify “doing the V* gesture very slowly and precisely”. Lastly, we show that our method can generalize to multiple people doing the motion while preserving the temporal consistency of the video. The Tune-A-Video baseline produces qualitatively worse results, as can be seen in Figure 5 and the supplementary video.
5 Conclusion
We present an approach for motion customization in text-to-video diffusion models. Our method can learn a new motion from a set of few exemplar videos of different subjects performing the same gesture. We conduct a thorough ablation study that identifies the key components of the method and evaluate them with respect to motion accuracy. We demonstrate qualitative results of our method, specifically in scenarios where our method generalizes the motion to unseen actors, multiple people performing the motion, and different viewpoints. Additionally, we show how our method can be extended to multi-modal motion and appearance customization to generate videos with custom subjects performing custom motions.
Supplementary Material
This supplemental contains the following material:
Accompanying webpage of qualitative results of our method, the videos shown in the main paper and comparison with the pre-trained text-to-video generations.
(Section A) We provide additional implementation details regarding our quantitative experiments.
(Section B) We discuss the tradeoff between the motion accuracy, appearance alignment and copying score across training iterations.
(Section C) We show qualitative examples along with their copying score.
(Section D) We show the original model’s (i.e. before customization) performance on the Jester dataset.
Appendix A Implementation Details
Appendix B Tradeoff between motion accuracy, appearance alignment, and copying score
When customizing a model with a new motion, we observe a natural bias-variance tradeoff. The longer we train the model, the better motion accuracy becomes, we memorize elements of the training examples, and the model becomes less diverse. We illustrate this in Figure 7. With more training iterations, the network memorizes the gesture as well as visual elements of the training examples. These findings are indicated by both the motion accuracy and copying score increasing. On the other hand, the CLIP appearance score, decreases. The CLIP appearance is defined as an average CLIP score between the part of the text prompt that describes a person’s appearance and the video frames (i.e. in a prompt “A female firefighter doing the V* sign”, the CLIP appearance score accounts for “A female firefighter”). This score is a proxy for visual diversity, since our test set contains visual description of a person, it measures how accurate the description is. As we increase the number of iterations, the videos become less diverse, hence the score decreases. This trend is persistent across the models from our ablation study. To find a balance between those factors, we perform early stopping for all models when the CLIP appearance score reaches 0.265 (i.e. when the image still reasonably well corresponds to the text description).
Appendix C Copying score
Copying score, as defined in 4.2 measures how much of the training data appearance leaks into the generated data. The score corresponds to the percentage of examples in the test set that are above a specified threshold of the SSCD score with any of the training examples. We find that the SSCD copy detection score [29] is a good proxy for determining the memorization of the training examples’ appearance. We measure the SSCD detection as a maximum score between any frame of the training videos and any video frame from the test set. We empirically determine that pairs of videos with a score greater than 0.25 share significant visual similarities. As seen in Figure 8, the similarities can be either the background or the appearance of the person performing the gesture.
Appendix D Pre-trained model’s performance on motions from the Jester dataset.
To test whether the gestures from the Jester dataset, are well-represented in the pre-trained text-to-video model, we evaluate the model with a similar experiment as our ablation experiment (4.2). We query the pre-trained text-to-video model with the prompts from the test set (Section A), replacing the “doing the sks gesture” with a natural language description of the class, e.g., for the class “Drumming Fingers” - “drumming finger”. We then classify the videos with the pre-trained network trained on the Jester dataset. No videos were correctly classified because (i) the gestures are not well represented in the model and (ii) the generated videos do not adhere to the spatial assumption of the Jester dataset (where a single person is performing a gesture in front of a camera).